### v1

# Supersymmetric scale-separated AdS3 vacua of type IIB

Abstract
I construct supersymmetric, parametrically scale-separated AdS3 vacua of type IIB string theory. These arise as compactifications with orientifold planes on specific seven-dimensional solvmanifolds admitting co-closed $G_{2}$-structures, preserving minimal supersymmetry. There are solutions that include either one set or four sets of intersecting O5-planes in the smeared approximation, and parametric scale separation can be achieved by tuning unbounded fluxes to infinity. Additionally, the putative holographic field theory operators that are dual to the lightest scalars in the gravitational theory have integer conformal dimensions at tree level, aligning with other scale-separated models of type II string theory.
preprint: UUITP-04/25
1 Introduction
Establishing full moduli stabilisation while also realising a large enough hierarchy between the cosmological radius and the Kaluza-Klein radius remains a tremendous challenge in flux compactifications. Nevertheless, this is one of the most standard features that phenomenological compactifications need to exhibit, such that the vacua are truly lower-dimensional. However, most AdS compactifications of string theory do not satisfy the scale separation condition:
$\frac{L_{\mathrm{KK}}}{L_{\mathrm{AdS}}}\ll 1\,,$ (1)
where $L_{\mathrm{KK}}$ and $L_{\mathrm{AdS}}$ are the KK and AdS radius respectively. Some recent progress has provided strong arguments against scale separation in specific settings. For instance, refs. Cribiori:2022trc ; Bobev:2023dwx ; Perlmutter:2024noo indicate that scale separation for compactifications of several theories with extended supersymmetry is unattainable. Moreover, it has been argued in ref. Gautason:2015tig that for compactifications without sources, scale separation cannot occur for isotropic manifolds of positive scalar curvature, more precisely for when the KK scale decouples from the scalar curvature, and due to the lack of concrete examples it was conjectured not to happen in ref. Collins:2022nux . All these findings support the strong version of the Anti-de Sitter distance conjecture Lust:2019zwm , implying that no supersymmetric scale-separated vacua exist. However, the minimally supersymmetric theories with sources are not constrained yet by the arguments of refs. Cribiori:2022trc ; Bobev:2023dwx ; Perlmutter:2024noo ; Gautason:2015tig ; Collins:2022nux , and indeed, there are a few scale-separated models in string theory, therefore violating the Anti-de Sitter distance conjecture Lust:2019zwm . Nevertheless, their validity is questioned and is an active research topic in the context of the swampland program. For a recent review on this issue, see ref. Coudarchet:2023mfs .
The current state of the art of supersymmetric scale-separated vacua can be summarised as a type IIA vs. IIB dichotomy: in type IIA, there are scale-separated vacua where all moduli are stabilised using fluxes and orientifold planes, hence they are referred to as classical vacua. The separation of scales is achieved by tuning unbounded, integer-quantised fluxes to infinity. Examples in four dimensions are compactifications on manifolds with SU(3)-structure, usually toroidal orbifolds, e.g. the DGKT solutions in massive IIA DeWolfe:2005uu ; Camara:2005dc and analogue solutions in massless IIA Cribiori:2021djm ; Carrasco:2023hta , of which some can be lifted to solutions of M-theory which appear to have weak $G_{2}$-structure VanHemelryck:2024bas . Examples in three dimensions are compactifications on manifolds with $G_{2}$-holonomy, as considered in Farakos:2020phe ; VanHemelryck:2022ynr . [^1] See also refs. Farakos:2023nms ; Farakos:2023wps ; Tringas:2023vzn for variations on these setups. Although the holographic dual field theories to these models are not known, refs. Conlon:2021cjk ; Apers:2022tfm ; Quirant:2022fpn found, quite surprisingly, that the mass spectrum of the scalars in the four-dimensional scale-separated vacua is such that the putative holographic dual operators in the field theory have integer conformal dimensions. [^2] This feature is expected to break down by small one-loop corrections, see ref. Plauschinn:2022ztd . This feature does not appear in the three-dimensional vacua of ref. Farakos:2020phe , as shown in ref. Apers:2022zjx .
However, these constructions have received criticism from several angles. For instance, these models all contain O6-planes that intersect in the covering space of the orbifold, and their charge density has been smeared over compact space, i.e. they are not treated as the codimension-3 objects that they are. However, this criticism has been refuted for a large part as the local orientifold backreaction has been computed and shown to be small Baines:2020dmu ; Junghans:2020acz ; Marchesano:2020qvg ; Andriot:2023fss ; Emelin:2024vug . Moreover, the O6-planes need not intersect once the orbifold singularities are blown up Junghans:2023yue . Nevertheless, the massive type IIA models cannot be lifted to M-theory due to the non-zero Romans mass. The models in massless type IIA Cribiori:2021djm ; Carrasco:2023hta overcome some of these issues, as they do have the potential to lift to M-theory on manifolds with weak $G_{2}$-structure, as recently shown in ref. VanHemelryck:2024bas and where the O6-planes geometrise. Additionally, it has been argued recently in ref. Montero:2024qtz that the massive IIA backgrounds are in tension with the weak gravity conjecture for branes Arkani-Hamed:2006emk , as the moduli space of the 3d theory living on a D4-domain wall is lifted due to the absence of a parity symmetry on the brane worldvolume. The massless type IIA vacua do not suffer from this problem.
The situation in type IIB is quite different. Non-perturbative effects are usually necessary for full moduli stabilisation and obtaining scale separation. Examples are four-dimensional Calabi-Yau compactifications with three-form fluxes and a gaugino condensate or Euclidean D3-instantons, such as in the KKLT and LVS scenarios Kachru:2003aw ; Conlon:2005ki ; Balasubramanian:2005zx . Recent progress has been made in finding solutions with a large hierarchy of scales (i.e. finding a small flux superpotential), see for example refs. Demirtas:2019sip ; Demirtas:2020ffz ; Demirtas:2021nlu ; Demirtas:2021ote ; McAllister:2024lnt , however, these scenarios have been criticised on different fronts as well, see e.g. refs. Bena:2020xrh ; Bena:2018fqc ; Gao:2020xqh ; Gao:2022fdi ; Gao:2022uop ; Lust:2022lfc for the most recent criticism. So far, no supersymmetric and scale-separated vacua of type IIB have been found where only classical ingredients stabilise all moduli, while maintaining weak coupling and large volume. [^3] The 4d vacua of refs. Petrini:2013ika ; Caviezel:2009tu are not well-controlled, as some one-cycles are becoming small in string units, see ref. Cribiori:2021djm .
However, this paper presents the first such examples. These are realised by compactifying type IIB string theory on seven-dimensional manifolds with a co-closed $G_{2}$-structure. Compactifications on such backgrounds have been considered before in refs. Emelin:2021gzx ; Passias:2020ubv ; VanHemelryck:2022ynr . Here, two specific orbifolds of a particular seven-dimensional solvmanifold are used. This solvmanifold has been described in the literature before, see e.g. refs. Andriot:2015sia ; Manero2020315 . After orientifolding these specific spaces with an O5-involution, one obtains one class of manifolds with four (intersecting) sets of (smeared) O5-planes and another class with only one set of O5-planes. The former is reminiscent of the type IIA vacua with 4 sets of (intersecting) O6-planes, whereas the latter case can be regarded as the supersymmetric cousin of a scale-separated, but non-supersymmetric type IIB vacuum found in ref. Arboleya:2024vnp , which are related through skew-whiffing. The analysis here uses the tools of refs. Emelin:2021gzx ; Passias:2020ubv ; VanHemelryck:2022ynr , but is different in the sense that in those references, models with seven (intersecting) sets of O5-planes were considered instead of fewer. Particularly in ref. Emelin:2021gzx , it was argued that scale separation can be achieved in their setup if one includes additional D5-branes to cancel some of the tadpoles, and if one can make the curvature small compared to the compactification radius. However, the latter clashes with geometric flux quantisation when the manifold is isotropic, and no further explicit models were considered either. The setups of this paper circumvent these issues and do not need the inclusion of D5-branes. Additionally, when computing the masses of the scalars in the untwisted sector, one obtains that the conformal dimensions of the corresponding operators in the putative holographic dual theory, are all integer, sharing this feature with their four-dimensional counterparts.
This paper is organised as follows: section 2 discusses the seven-dimensional solvmanifold and the different orbifold groups that it is quotiented by, such that an orientifold involution produces the right O5-plane content for the type IIB solutions of interest. Section 3 contains a brief review of type IIB compactifications on manifolds with $G_{2}$-structure, and discusses the specific compactifications on the manifolds introduced in the previous section and their scale separation feature. In section 4, the mass spectrum and conformal dimensions are discussed. Finally, the results are summarised in section 5.
2 A seven-dimensional solvmanifold, orbifolded and orientifolded
Before considering the details of three-dimensional flux compactifications of type IIB string theory, it is useful to focus on the specific compact space that is used later in section 3. In particular, solvmanifolds are considered throughout this paper, which are group manifolds arising from a solvable group. There exists a variety of such solvmanifolds across different dimensions, amongst which many have been used for compactifications before, for instance in refs. Grana:2006kf ; Andriot:2010ju ; Danielsson:2011au ; Grana:2013ila ; Andriot:2015sia ; Andriot:2022way ; Andriot:2022yyj . Here, a very specific seven-dimensional solvmanifold is considered. It arises as a twisted torus, denoted by $\tilde{\mathbb{T}}^{7}$, for which the metric in the vielbein parametrisation is
$\mathrm{d}s_{7}^{2}=\sum_{i=1}^{7}L_{i}^{2}(e^{i})^{2}\,,$ (2)
where the $e^{i}=e^{i}_{m}\mathrm{d}x^{m}$ are the unit length local frame one-forms and the $L_{i}$ are the associated radii, expressed in string units ($2\pi\ell_{s}=1)$. The particular example of a seven-dimensional solvmanifold that is of interest here, is realised by the following Mauer-Cartan equations:
$\begin{split}&\mathrm{d}e^{1}=\omega_{a}\;e^{2}\wedge e^{7},\qquad\mathrm{d}e^{2}=-\;\omega_{a}e^{1}\wedge e^{7}\\ &\mathrm{d}e^{3}=\omega_{b}\;e^{4}\wedge e^{7},\qquad\mathrm{d}e^{4}=-\;\omega_{b}e^{3}\wedge e^{7}\\ &\mathrm{d}e^{5}=\omega_{c}\;e^{6}\wedge e^{7},\qquad\mathrm{d}e^{6}=-\;\omega_{c}e^{5}\wedge e^{7}\,.\end{split}$ (3)
These Mauer-Cartan equations can be integrated into the following expressions for the one-forms, in terms of the coordinates $x^{1},...,x^{7}$:
$\begin{split}&e^{1}=\cos\left(\omega_{a}x^{7}\right)\mathrm{d}x^{1}-\sin\left(\omega_{a}x^{7}\right)\mathrm{d}x^{2}\\ &e^{2}=\sin\left(\omega_{a}x^{7}\right)\mathrm{d}x^{1}+\cos\left(\omega_{a}x^{7}\right)\mathrm{d}x^{2}\\ &e^{7}=\mathrm{d}x^{7}\,.\end{split}$ (4)
and for the other directions idem upon the simultaneous cyclic permutation of (1,3,5), (2,4,6) and $(\omega_{a},\omega_{b},\omega_{c})$. This manifold has been considered before, for example in refs. Andriot:2015sia ; Manero2020315 . In essence, this seven-dimensional solvmanifold can be seen as three different copies of the 3d solvmanifold $E_{2}$ (also called ISO(2)) with one common direction (the seventh), see e.g. refs. Grana:2006kf ; Grana:2013ila ; Andriot:2015sia . The three-dimensional space can be divided by different lattices to make it compact, each lattice leading to a different cohomology, see refs. Grana:2013ila ; Andriot:2015sia . However, for this paper, it is enough to take the standard torus lattice $x^{i}\simeq x^{i}+1$, as the Betti numbers will be reduced through an orbifold. Consequentially, this means that the structure constants $\omega_{a,b,c}$ have to be quantised by $2\pi\mathbb{Z}$ when all the $e^{i}$’s have to be preserved by the lattice.
The orbifolds and orientifold
To project out certain forms of the seven-dimensional manifold, and to generate multiple orientifold plane images, it is useful to consider orbifolds of this space. A nice feature of the solvmanifold is that many of the orbifold actions compatible with the flat seven-torus $\mathbb{T}^{7}$, are also compatible here. For example, the standard $\Gamma=\mathbb{Z}_{2}^{3}$ orbifold of the flat torus $\mathbb{T}^{7}$ can be considered here as well. The elements of $\Gamma$ can be written as follows:
$\begin{split}&\alpha(x)=(-x^{1},-x^{2},x^{3},x^{4},-x^{5},-x^{6},x^{7})\\ &\beta(x)=(x^{1},x^{2},-x^{3},-x^{4},-x^{5},-x^{6},x^{7})\\ &\gamma(x)=(-x^{1},x^{2},-x^{3},x^{4},-x^{5},x^{6},-x^{7})\,.\end{split}$ (5)
These three $\mathbb{Z}_{2}$ actions preserve the Mauer-Cartan equations (3), square to one and each have 16 fixed points, just as their combinations $\alpha\beta$, $\gamma\beta$, $\alpha\gamma$, $\alpha\beta\gamma$. For example, the fixed points of $\gamma(x)=x$ are at $x^{1,3,5,7}=\{0,\frac{1}{2}\}$, for all $x^{2,4,6}$. After orbifolding the twisted torus, the space should be orientifolded by an O5-involution for further purposes of this paper, similar to ref. Emelin:2021gzx . Such an involution that also leaves the Mauer-Cartan equations (3) invariant, is the following:
$\sigma(x)=(-x^{1},x^{2},-x^{3},x^{4},-x^{5},x^{6},-x^{7})\,,$ (6)
acting on the coordinates in the same way as $\gamma$. The hypersurfaces of fixed points of this orientifold involution correspond to O5-planes, as they are of codimension 4. Since the involution acts the same as $\gamma$, the orientifold $\sigma$ has the same fixed points as $\gamma$, at $x^{1,3,5,7}=\{0,\frac{1}{2}\}$. Moreover, all orbifold images of $\sigma$, i.e. $\sigma\alpha$, $\sigma{\alpha\beta}$, … have fixed points too, which all lead to different sets of O5-planes. Finally, $\sigma\gamma$ leaves everything invariant and leads to an O9-plane, meaning that one can consider type I string theory as well, after including 32 D9-branes. This orbifold and orientifold were already considered in ref. Emelin:2021gzx , and the inclusion of the O9-plane was useful for matching results with 3d compactifications of heterotic string theory, see e.g. deIaOssa:2019cci ; delaOssa:2021cgd ; delaOssa:2024dzo , due to heterotic - type I string theory duality. Nevertheless, it is not well-understood how the orbifold singularities of such spaces can be resolved.
For this work, it is important to reduce the number of O5-plane sets, and additionally remove the O9-plane so that one remains within type IIB string theory. This can be achieved by changing the orbifold actions by introducing shifts as well. Such types of orbifolds were crucial in Joyce’s construction of compact spaces with $G_{2}$-holonomy Joyce:1996i ; Joyce:1996ii . As a first example, one can consider dividing out by $\hat{\Gamma}=\mathbb{Z}_{2}^{3}$, with the same action $\alpha$, but where shifts are introduced in $\beta$ and $\gamma$:
$\begin{split}&\hat{\alpha}(x)=\biggl{(}-x^{1},-x^{2},x^{3},x^{4},-x^{5},-x^{6},x^{7}\biggr{)}\\ &\hat{\beta}(x)=\biggl{(}x^{1},x^{2},-x^{3},-x^{4},-x^{5},-x^{6},x^{7}+\frac{1}{2}\biggr{)}\\ &\hat{\gamma}(x)=\biggl{(}-x^{1}+\frac{1}{2},x^{2},-x^{3}+\frac{1}{2},x^{4},-x^{5}+\frac{1}{2},x^{6},-x^{7}\biggr{)}\,.\end{split}$ (7)
Only the elements $\hat{\alpha}$ and $\hat{\gamma}$ have fixed points (16 each), all the other combinations do not. Notice that the extra shift in $x^{7}$ requires one to quantise the structure constants $\omega_{a,b,c}$ in $4\pi\mathbb{Z}$ instead of $2\pi\mathbb{Z}$, if one wants all one-forms $e^{i}\to\pm e^{i}$ to just flip sign under all orbifold actions. [^4] Note that a similar quantisation condition also appeared for nilmanifolds, e.g. in ref. Cribiori:2021djm . According to the criteria of Joyce Joyce:1996i ; Joyce:1996ii , this orbifold on the flat $\mathbb{T}^{7}$ has singularities of the ‘good’ type, meaning that they do not intersect and that a blow-up procedure can resolve them. Every fixed point is locally a three-torus on the flat $\mathbb{T}^{7}$. The group element $\hat{\gamma}$ reduces the set of independent fixed $\mathbb{T}^{3}$’s of $\hat{\alpha}$ by half, to 8, hence there are eight singularities of the local form $\mathbb{R}^{4}/\mathbb{Z}_{2}\times\mathbb{T}^{3}$. Similarly, the fixed points of $\hat{\gamma}$ are shuffled around by both $\hat{\alpha}$ and $\hat{\beta}$, leading to only 4 independent singularities. The blow-up procedure generates three local three-cycles for every independent singularity, leading to $(8+4)\times 3=36$ additional three-cycles, whose associated blow-up moduli are said to belong to the twisted sector. There are also 12 local two-cycles from the blow-up. Although it is not straightforward to extrapolate this to the twisted torus $\tilde{\mathbb{T}}^{7}$ here, one could argue that, in a neighbourhood of the fixed points of $\hat{\gamma}$, the twisted torus looks locally like the flat one. For $\hat{\alpha}$ on the other hand, the fixed loci are the three-dimensional solvmanifolds $E_{2}$ instead of $\mathbb{T}^{3}$’s. However, if one would choose to set $\omega_{b}=0$, then they are $\mathbb{T}^{3}$’s, and the arguments above might apply again, although one has to treat this reasoning with caution.
Further orientifolding with the same involution $\sigma$ of eq. (6) results in fixed points of codimension four for $\sigma$, $\sigma\hat{\alpha}$, $\sigma\hat{\beta}$ and $\sigma{\alpha\hat{\beta}}$ , leading to 4 sets of O5-planes. These are summarised in Table 1. The actions $\sigma\hat{\gamma}$, $\sigma{\hat{\alpha}\hat{\gamma}}$, $\sigma{\hat{\beta}\hat{\gamma}}$, $\sigma{\hat{\alpha}\hat{\beta}\hat{\gamma}}$ have no fixed points: for each of them, there is at least one coordinate $x^{i}$ that is mapped into $x^{i}+1/2$, hence no further orientifold images are created. The orientifold involution does not reduce the amount of independent orbifold singularities and therefore does not seem to affect the twisted sector.
$x^{1}$ $x^{2}$ $x^{3}$ $x^{4}$ $x^{5}$ $x^{6}$ $x^{7}$
O5σ -- $\times$ -- $\times$ -- $\times$ --
[…]
Table 1: The different orientifold planes after orientifolding the orbifold $\hat{\Gamma}$. The crosses indicate the directions parallel to the O5-plane, while the dashes correspond to the directions in which the O5-planes are localised, at the values $\{0,1/2\}$ and $\{1/4,3/4\}$, the latter being indicated with an asterisk.
Another example arises from introducing more shifts, for instance by quotienting with the $\check{\Gamma}=\mathbb{Z}_{2}^{3}$ group with the following group elements:
$\begin{split}&\check{\alpha}(x)=\left(-x^{1},-x^{2}+\frac{1}{2},x^{3},x^{4}+\frac{1}{2},-x^{5},-x^{6},x^{7}\right)\\ &\check{\beta}(x)=\left(x^{1},x^{2}+\frac{1}{2},-x^{3},-x^{4},-x^{5},-x^{6}+\frac{1}{2},x^{7}\right)\\ &\check{\gamma}(x)=\left(-x^{1}+\frac{1}{2},x^{2},-x^{3}+\frac{1}{2},x^{4},-x^{5}+\frac{1}{2},x^{6},-x^{7}\right)\,.\end{split}$ (8)
An orbifold like this appeared in ref. Andriolo:2018yrz but without the shifts in $\check{\gamma}$. Only the orbifold action $\check{\gamma}$ has fixed points, and again, Joyce’s criterion guarantees that this orbifold of the flat torus has singularities which can be blown up, of which there are only 4. Furthermore, only the orientifold involution $\sigma$ has fixed points, all orbifold images do not have any. Hence there is only one set of O5-planes. The orientifold involution does not further identify different orbifold singularities here either.
$x^{1}$ $x^{2}$ $x^{3}$ $x^{4}$ $x^{5}$ $x^{6}$ $x^{7}$
O5σ $-^{*}$$\times$ $-^{*}$$\times$ $-^{*}$$\times$ --
Table 2: The different orientifold planes after orientifolding the orbifold $\check{\Gamma}$. The crosses indicate the directions parallel to the O5-plane, while the dashes correspond to the directions in which the O5-planes are localised, at the values $\{0,1/2\}$ and $\{1/4,3/4\}$, the latter being indicated with an asterisk.
When the solvmanifold is quotiented by one of the orbifold groups $\Gamma$, $\hat{\Gamma}$ or $\check{\Gamma}$, the untwisted Betti numbers are reduced. Indeed, there are no one- or two-forms that are preserved under the orbifold, i.e. $b_{1}=b_{2}=0$ and there are only seven three-forms and seven four-forms that are preserved, i.e. $b_{3}=7$. A basis for them is given by
$\begin{split}&\Phi^{i}=\left\{e^{127},\,e^{347},\,e^{567},\,e^{136},\,e^{235},\,e^{145},\,-e^{246}\right\}\,,\\ &\Psi^{i}=\left\{e^{3456},\,e^{1256},\,e^{1234},\,e^{2457},\,e^{1467},\,e^{2367},\,-e^{1357}\right\}\,.\end{split}$ (9)
where the short-hand notation $e^{a...b}=e^{a}\wedge...\wedge e^{b}$ is used. Note that all seven four-forms are closed under the exterior derivative by virtue of the Mauer-Cartan equations (3). The same is true for the first three four-forms, the other four are not closed. Consequentially, the seven-dimensional manifold admits a $G_{2}$-structure with a left-invariant three-form $\Phi$ and its associated four-form $\star\Phi$, which can be parametrised as follows:
$\begin{split}\Phi&=\;s_{127}\>e^{127}+s_{347}\>e^{347}+s_{567}\>e^{567}+s_{136}\>e^{136}+s_{235}\>e^{235}+s_{145}\>e^{145}-s_{246}\>e^{246}\\ \star\Phi&=\;s_{3456}\>e^{3456}+s_{1256}\>e^{1256}+s_{1234}\>e^{1234}+s_{2457}\>e^{2457}+s_{1467}\>e^{1467}\\ &\quad+s_{2357}\>e^{2367}-s_{1357}\>e^{1357}\,,\end{split}$ (10)
where $s_{ijk}=L_{i}L_{j}L_{k}$ and $s_{ijkl}=L_{i}L_{j}L_{k}L_{l}$. Again, the forms $\Phi$ and $\star\Phi$ are invariant under all orbifold actions $\Gamma$, $\hat{\Gamma}$ and $\check{\Gamma}$, and they are even under the orientifold involution $\sigma$ as well. The $G_{2}$ left-invariant three- and four-form do generally not need to be closed. Their non-closure can be described by the torsion classes, classified in different irreducible representations of the $G_{2}$ group:
$\mathrm{d}\Phi=W_{1}\star\Phi+W_{7}\wedge\Phi+W_{27}\,,\qquad\mathrm{d}\star\Phi=\frac{4}{3}W_{7}\wedge\star\Phi+W_{14}\wedge\Phi.$ (11)
The Ricci curvature can also be expressed in terms of the torsion classes:
$R_{7}=\frac{21}{8}|W_{1}|^{2}+\frac{10}{3}|W_{7}|^{2}+4\star\mathrm{d}\star W_{7}-\frac{1}{2}|W_{14}|^{2}-\frac{1}{2}|W_{27}|^{2}\,.$ (12)
The Mauer-Cartan relations (3) imply that the 7d solvmanifolds admit a co-closed $G_{2}$-structure, but it does not necessarily have to be closed:
$\mathrm{d}\Phi=W_{1}\star\Phi+W_{27}\,,\qquad\mathrm{d}\star\Phi=0\,.$ (13)
However, by choosing the structure constants appropriately, i.e. $\omega_{\Sigma}=\omega_{a}+\omega_{b}+\omega_{c}=0$ and by taking $L_{1}=L_{2}$, $L_{3}=L_{4}$ and $L_{5}=L_{6}$, the solvmanifold has a closed three-form as well, which was already noticed in refs. Andriot:2015sia ; Manero2020315 . Nevertheless, co-closed $G_{2}$-structures play an important role in minimally supersymmetric type IIB compactifications to AdS3, as is discussed below.
3 Scale-separated solutions of type IIB with co-closed $G_{2}$-structures
This section uses the seven-dimensional solvmanifolds introduced above to construct AdS3 solutions of type IIB string theory. As they admit a co-closed $G_{2}$-structure, the tools of refs. Emelin:2021gzx ; VanHemelryck:2022ynr ; Passias:2020ubv can be used immediately. Compactifiying on such spaces preserves only 1/8 of the supersymmetries, and orientifolding breaks another half of them, resulting in minimal supersymmetry in three dimensions. The ten-dimensional metric can be written as a direct sum as follows:
$\mathrm{d}s_{10}^{2}=\mathrm{e}^{2A}\mathrm{d}s_{3}^{2}+\mathrm{d}s_{7}^{2}\,,$ (14)
where $A$ is the warp factor, which is a function on the seven-dimensional space, governed by the metric $\mathrm{d}s_{7}^{2}$ which measures distances in string units. To find supersymmetric solutions, one can solve the killing spinor equations of type IIB string theory and impose the Bianchi identities. The killing spinor equations were rewritten in ref. Dibitetto:2018ftj in terms of bispinors for general backgrounds and applied in VanHemelryck:2022ynr ; Passias:2020ubv for $G_{2}$-structure backgrounds in type IIB. They require that the $G_{2}$-structure must be co-closed, a combination of the warping and the dilaton must be constant and that the NSNS three-form must vanish:
$\mathrm{d}\star\Phi=0\,,\qquad H=0\,,\qquad\mathrm{e}^{2A-\phi}=g_{s}^{-1}\,.$ (15)
Furthermore, the geometry is fixed by the RR fluxes, however, it is more convenient to take the other perspective and express the fluxes in terms of the geometry
$F_{1}=0$ (16)
$F_{3}=-g_{s}^{-1}\mathrm{e}^{-3A}\star\mathrm{d}\left(\mathrm{e}^{A}\Phi\right)+2\mu g_{s}^{-1}\mathrm{e}^{-3A}\Phi$ (17)
$F_{5}=0$ (18)
$F_{7}=-2\mu g_{s}^{-1}\;\text{vol}_{7}.$ (19)
The constant $\mu$ is related to the inverse AdS radius, i.e. $\mu^{2}=1/L_{\mathrm{AdS}}^{2}$. Finally, the torsion class $W_{1}$ and AdS scale are also related by the killing spinor equations:
$W_{1}=\frac{12}{7}\mu\>\mathrm{e}^{-A}.$ (20)
The RR three-form flux must also satisfy its Bianchi identity, which can be sourced by D5-branes or O5-planes. Here, it is enough to consider O5-planes only, and the relevant Bianchi identity in string units becomes:
$\mathrm{d}F_{3}=-\sum_{i,j}\delta(\Sigma_{\text{O5}_{\perp,i,j}})\,.$ (21)
The sum over $i$ runs over distinct O5-plane sets in different directions, while the sum over $j$ accounts for multiple O5-plane images parallel to those directions. It is useful to work in the smeared approximation for the O-planes, where the Dirac-delta forms are replaced by the unit volume forms of the cycles that the O5’s are orthogonal to, i.e. $\sum_{j}\delta(\Sigma_{\text{O5}_{\perp,i,j}})\to j_{\text{O5}_{i}}=\rho_{\mathrm{O5}}\text{vol}(\Sigma_{\text{O5}_{\perp,i}})$ where unit volume is understood here and $\rho_{\mathrm{O5}}$ is a constant charge density. In that case, the warping and dilaton profile are constant, and the solution for $F_{3}$ reduces to
$g_{s}F_{3}=\frac{W_{1}}{6}\Phi-\star W_{27}\,.$ (22)
One concludes that, for any co-closed $G_{2}$-structure manifold, there exists a supersymmetric 3d compactification, provided that the manifold has the proper orientifold content such that the $F_{3}$-Bianchi identity is satisfied (in the smeared approximation).
As stated above, the different orbifolds $\hat{\Gamma}$ and $\check{\Gamma}$ of the seven-dimensional solvmanifold are examples of co-closed $G_{2}$-structures and can therefore be used for these type IIB compactifications. These are considered below.
Solutions on $\tilde{\mathbb{T}}^{7}/\hat{\Gamma}$
First, consider the space $\tilde{\mathbb{T}}^{7}/\hat{\Gamma}$, orbifolded by $\hat{\Gamma}$ of section 2 which, upon orientifolding, admits four sets of orientifold planes as in Table 1. It turns out to be useful to look at the situation where $L_{1}=L_{2}$, $L_{3}=L_{4}$ and $L_{5}=L_{6}$, which simplifies the setup but such that not all three-cycle volumes are the same. For this choice, the exterior derivative on $\Phi$ gives the simple expression:
$\mathrm{d}\Phi=L_{1}L_{3}L_{5}\omega_{\Sigma}\left(e^{2457}+e^{1467}+e^{2367}-e^{1357}\right),\quad\omega_{\Sigma}=\omega_{a}+\omega_{b}+\omega_{c}\,,\quad W_{1}=\frac{4\omega_{\Sigma}}{7L_{7}}\,.$ (23)
It is important to note that when $\omega_{\Sigma}$ vanishes, the $G_{2}$-structure is both closed and co-closed. This was already noticed in refs. Andriot:2015sia ; Manero2020315 . In that case, $\mu$ also vanishes due to eq. (20) and the external spacetime is Minkowski instead of Anti-de Sitter. So for now, one can take $\omega_{\Sigma}\neq 0$. Even in this case, the Ricci curvature $R_{7}$ vanishes:
$R_{7}=\frac{21}{8}|W_{1}|^{2}-\frac{1}{2}|W_{27}|^{2}=0.$ (24)
Furthermore, the supersymmetric solutions above, eqs. (17)-(19) determine $F_{3}$ and $F_{7}$ completely in terms of the geometry:
$F_{3}=\left(f_{3,1}\;e^{127}+f_{3,3}\;e^{347}+f_{3,5}\;e^{567}\right)-\tilde{f}_{3}\left(e^{136}+e^{235}+e^{145}-e^{246}\right),\quad F_{7}=-f_{7}\,e^{1234567},$ (25)
$f_{3,i}=\frac{2L_{i}^{2}\omega_{\Sigma}}{3g_{s}}\,,\qquad\tilde{f}_{3}=\frac{L_{1}L_{3}L_{5}\omega_{\Sigma}}{3g_{s}L_{7}}\,,\qquad f_{7}=\frac{2\omega_{\Sigma}(L_{1}L_{3}L_{5})^{2}}{3g_{s}}\,,$ (26)
where the index $i$ runs over $\{1,3,5\}$. The radii and string coupling can also be expressed in terms of the flux parameters by inverting the last identities:
$L_{i}=\left(\frac{f_{7}\;f_{3,i}^{2}}{f_{3,1}\;f_{3,3}\;f_{3,5}}\right)^{\frac{1}{4}},\quad L_{7}=\frac{\left(f_{7}\;f_{3,1}\;f_{3,3}\;f_{3,5}\right)^{\frac{1}{4}}}{2\tilde{f}_{3}},\quad g_{s}=\frac{2\omega_{\Sigma}}{3}\left(\frac{f_{7}}{f_{3,1}\;f_{3,3}\;f_{3,5}}\right)^{\frac{1}{2}}\,.$ (27)
Finally, $F_{3}$ is not closed, but due to the Mauer-Cartan equations (3), only the components including $\tilde{f}_{3}$ are responsible for that:
$\mathrm{d}F_{3}=-\tilde{f}_{3}\omega_{\Sigma}\left(e^{2457}+e^{1467}+e^{2367}-e^{1357}\right)\,,$ (28)
and since $\tilde{f}_{3}\omega_{\Sigma}>0$, the Bianchi identity of $F_{3}$ (21) should be solved by four sets of O5-planes. And indeed, as discussed in the previous sections, by orientifolding $\tilde{\mathbb{T}}^{7}/\hat{\Gamma}$, one gets, upon smearing, precisely the O5-planes required as in Table 1. Note also that the components of $F_{3}$ including the $f_{3,i}$’s are closed and do not participate in the Bianchi identity. This implies that all the $f_{3,i}$’s are unbounded flux quanta that can be chosen to be arbitrarily large.
Solutions on $\tilde{\mathbb{T}}^{7}/\check{\Gamma}$
Another interesting option is the scenario on $\tilde{\mathbb{T}}^{7}/\check{\Gamma}$, where the radii are also related in pairs by $L_{2}=\sqrt{2}L_{1}$, $L_{4}=\sqrt{2}L_{3}$ and $L_{6}=\sqrt{2}L_{5}$. Additionally, it is interesting to choose all structure constants to be the same in size, but where one of them has the opposite sign, which can be taken to be $\omega_{a}$ without loss of generality, i.e. $-\omega_{a}=\omega_{b}=\omega_{c}$. For this choice, the non-closure of $\Phi$ is captured by:
$\mathrm{d}\Phi=-\sqrt{2}L_{1}L_{3}L_{5}\omega_{a}\left(2e^{2367}+2e^{2457}-e^{1357}\right),\qquad W_{1}=-\frac{3\sqrt{2}\omega_{a}}{7L_{7}}\,,$ (29)
and the scalar curvature does not vanish. Using this space, the RR three-form flux is of a similar form as above:
$F_{3}=-\left(f_{3,1}\;e^{127}+f_{3,3}\;e^{347}+f_{3,5}\;e^{567}\right)+\tilde{f}_{3}\left(e^{136}-e^{235}+e^{145}-2e^{246}\right),\quad F_{7}=f_{7}\,e^{1234567},$ (30)
$f_{3,i}=\frac{L_{i}^{2}\omega_{a}}{g_{s}}\,,\qquad\tilde{f}_{3}=\frac{L_{1}L_{3}L_{5}\omega_{a}}{g_{s}L_{7}}\,,\qquad f_{7}=\frac{2\omega_{a}(L_{1}L_{3}L_{5})^{2}}{g_{s}}\,,$ (31)
where the index $i$ runs again through $\{1,3,5\}$. Expressing the scalars in terms of the fluxes gives:
$L_{i}=\left(\frac{f_{7}\;f_{3,i}^{2}}{2f_{3,1}\;f_{3,3}\;f_{3,5}}\right)^{\frac{1}{4}},\quad L_{7}=\frac{\left(f_{7}\;f_{3,1}\;f_{3,3}\;f_{3,5}\right)^{\frac{1}{4}}}{2^{\frac{1}{4}}\tilde{f}_{3}},\quad g_{s}=\omega_{a}\left(\frac{f_{7}}{2f_{3,1}\;f_{3,3}\;f_{3,5}}\right)^{\frac{1}{2}}\,.$ (32)
Note that there are some similarities but also key differences with the solution (25) on the previous orbifold. For example, the signs are different, especially the one for the 235-component, and the 246-component has an additional factor of 2. Nevertheless, the dependence on the flux parameters for the radii and string coupling is the same as what was considered for $\tilde{\mathbb{T}}^{7}/\hat{\Gamma}$. The RR three-form is not closed here either, and again, only the components including $\tilde{f}_{3}$ are responsible for that, only this time they generate just one component:
$\mathrm{d}F_{3}=-3\tilde{f}_{3}\,\omega_{a}\,\left(-e^{1357}\right)\,.$ (33)
To satisfy the Bianchi identity, only one set of (smeared) O5-planes is required. As indicated above, orientifolding $\tilde{\mathbb{T}}^{7}/\check{\Gamma}$ generates only one set of O5-planes along the proper directions, as in Table 2. Again, the $f_{3,i}$’s are not restricted by the tadpole and can be chosen to be arbitrarily large. Note that solutions on this space should be seen as the supersymmetric cousins of some of the solutions found in ref. Arboleya:2024vnp , which also include just one set of O5-planes and are non-supersymmetric. However, that paper uses another orbifold than the specific $\tilde{\mathbb{T}}^{7}/\check{\Gamma}$ and therefore might need additional D5-branes to cancel O5-charge along other directions.
Scale separation
Now it is time to investigate whether there is a separation of scales between the AdS scale and the volume of the three-cycles of the manifold. For this purpose, it is quintessential to make use of the fluxes that are unrestricted by tadpoles, which are parametrised here by the $f_{3,i}$’s and $f_{7}$. The flux parameter $\tilde{f}_{3}$ and structure constants $\omega_{a,b,c}$ are constrained by the $F_{3}$ Bianchi identity, as they have to be balanced against the O5-plane charge density. The latter is fixed, hence it cannot be tuned to infinity. Indeed, for both models, the Bianchi identity imposes that the product $\tilde{f}_{3}\omega_{\Sigma}$ for $\tilde{\mathbb{T}}^{7}/\hat{\Gamma}$ or $\tilde{f}_{3}\omega_{a}$ for $\tilde{\mathbb{T}}^{7}/\check{\Gamma}$ should be held fixed. However, all flux parameters are quantised, also $\tilde{f}_{3}$ and $\omega_{\Sigma}$ or $\omega_{a}$, which means that they should be held fixed separately: neither one of them can be made arbitrarily small at the expense of the other one becoming large. Another important remark is that the AdS scale is of similar size as the radius $L_{7}$, as
$L_{\mathrm{AdS}}^{-1}=|\mu|=\frac{7}{12}|W_{1}|\sim\omega_{\Sigma}L_{7}^{-1}\text{ or }\omega_{a}L_{7}^{-1}\,.$ (34)
However, this is not a problem, as there are no cycles whose volumes are determined by $L_{7}$ alone. This is one of the reasons why the twisted seven-torus had to be orbifolded such that the one-forms, especially $e^{7}$, are projected out. The volumes of the generalised three-cycles are then set by $L_{i}^{2}L_{7}$ and $L_{1}L_{3}L_{5}$ for $i\in\{1,3,5\}$, which satisfy
$\frac{L_{i}^{2}L_{7}}{L_{\mathrm{AdS}}^{3}}\sim\frac{f_{3,i}}{f_{3,1}\;f_{3,3}\;f_{3,5}}\,,\qquad\frac{L_{1}L_{3}L_{5}}{L_{\mathrm{AdS}}^{3}}\sim\frac{1}{f_{3,1}\;f_{3,3}\;f_{3,5}}\,,$ (35)
and indeed, by sending the $f_{3,i}$’s to infinity, one obtains parametric scale separation. To ensure that all cycles are large in string units and that the string coupling is small, one should take the $f_{3,i}$’s and $f_{7}$ large, such that
$\frac{f_{3,1}\;f_{3,3}\;f_{3,5}}{f_{3,\alpha}^{2}}\ll f_{7}\ll f_{3,1}\;f_{3,3}\;f_{3,5}\,.$ (36)
where $f_{3,\alpha}$ is the smallest flux amongst the $f_{3,i}$’s. Hence scale separation can be achieved in the controlled regime, i.e. at small string coupling and large volume. Finally, it is also interesting to see how the 3d Planck mass scales with the fluxes when expressed in string units. In 3d, this is given by
$M_{\text{Pl}}=g_{s}^{-2}\mathcal{V}_{7}\sim(f_{7}\;f_{3,1}\;f_{3,3}\;f_{3,5})^{\frac{3}{4}}\,.$ (37)
Here $\mathcal{V}_{7}=L_{1}L_{2}L_{3}L_{4}L_{5}L_{6}L_{7}$ is the volume of the compact space. This would lead in a putative holographic dual theory to a central charge that satisfies:
$c=\frac{2}{3}M_{\text{Pl}}L_{\mathrm{AdS}}\sim f_{7}\;f_{3,1}\;f_{3,3}\;f_{3,5}\implies f_{3,\alpha}^{-2}(f_{3,1}\;f_{3,3}\;f_{3,5})^{2}\ll c\ll(f_{3,1}\;f_{3,3}\;f_{3,5})^{2}\,.$ (38)
This would result in a large number of degrees of freedom in the putative holographic dual theory. Additionally, there would be a parametrically large gap in the spectrum. Such conformal field theories with this feature have not been found yet.
Comments on using weak $G_{2}$-structures
So far, co-closed $G_{2}$-structures have been considered with non-vanishing torsion class $W_{27}$. This torsion class allows for non-positive scalar curvature, which requires O5-planes to solve Bianchi identities. However, it is also worthwhile to contemplate setups for which $W_{27}$ vanishes. In that case, only $W_{1}$ survives and the manifold is said to have a weak $G_{2}$-structure. Notice that $W_{1}$ is related to the inverse AdS radius and hence taking it to be vanishing would lead to Minkowski setups instead. So consider a type IIB background for which $W_{1}\neq 0$ and $W_{27}=0$. It follows from eq. (22) that
$\mathrm{d}F_{3}=g_{s}^{-1}\frac{W_{1}^{2}}{6}\star\Phi\,.$ (39)
Consequentially, the Bianchi identity should be solved by an appropriate amount of D5-branes wrapping three-cycles. The flux is not restricted in that case, as the number of D5-branes can be chosen at will. However, if one wants scale separation, then one would need an anisotropic manifold for which the torsion class $W_{1}$ and hence the scalar curvature radius decouple from the KK radius, estimated by the three- and four-cycle volume scales. Indeed, if the manifold is isotropic, then $\mu\sim W_{1}\sim 1/\mathcal{V}_{7}^{1/7}$, leading to no separation of scales. Furthermore, all homogeneous and smooth weak $G_{2}$-manifolds have been classified, and they do not allow for this property. Nevertheless, a weak $G_{2}$-structure manifold with these properties has recently been proposed in ref. VanHemelryck:2024bas , arising from a lift of a massless type IIA solution Cribiori:2021djm to M-theory, and could therefore lead to a hierarchy of scales in these 3d compactifications as well. However, exploring this scenario in full detail requires a more complete understanding of this space, which is left for future research.
4 Masses and conformal dimensions
In refs. Conlon:2021cjk ; Apers:2022tfm ; Quirant:2022fpn , it was noticed that in the parametrically scale-separated AdS4 vacua of massive type IIA DeWolfe:2005uu ; Camara:2005dc , the masses of the scalars are such that their corresponding operators in the putative holographic field theory have integer conformal dimensions. The same feature arises in the scale-separated vacua of ref. Cribiori:2021djm , as they are morally T-dual to the vacua of refs.DeWolfe:2005uu ; Camara:2005dc , although on a different orbifold. In the three-dimensional vacua of ref. Farakos:2020phe , the situation is quite different, as it was shown in ref. Apers:2022zjx that the conformal dimensions are not integer there. Hence, it is worthwhile to investigate whether this feature also appears for the vacua constructed here.
As was explained well in ref. Emelin:2021gzx , there are no axions in the type IIB setup here, either because homology does not allow for it (e.g. there is no $B$ or $C_{2}$ axion as there are no two-forms in the untwisted sector) or because the orientifold projects them out (e.g. $C_{4}$ is odd under an O5-involution but there are no odd four-cycles in the manifold). So the only scalar fields that enter the potential are the seven three-cylce volumes (or equivalently, the seven radii) and the dilaton. Since the radii $L_{k}$ defined above parametrise volumes in string units, it is convenient to trade the 10d dilaton $\phi$ for the 3d dilaton $\delta$. Additionally, it is useful here to trade the three-cycle volume scalars for the radii, such that the field space metric is diagonal:
$g_{ij}\partial\sigma^{i}\partial\sigma^{j}=4\frac{(\partial\delta)^{2}}{\delta^{2}}+\sum_{k=1}^{7}\frac{(\partial L_{k})^{2}}{L_{k}^{2}}\,,$ (40)
where the relation between both dilatons is that
$\delta^{2}=\mathrm{e}^{-2\phi}L_{1}L_{2}L_{3}L_{4}L_{5}L_{6}L_{7}\,.$ (41)
The scalar potential was already obtained from dimensional reduction in type IIB in ref. Emelin:2021gzx and is given by: [^5] Note that there is a difference of a factor 4 for the first term of the potential when comparing with ref. Emelin:2021gzx , which is a result of normalising the kinetic terms differently.
$V=4\left(g^{ij}(\partial_{i}P)(\partial_{j}P)-P^{2}\right)\,,$ (42)
where $g_{ij}$ is the field space metric and $P$ is the real superpotential first defined in ref. Emelin:2021gzx and derived in terms of bispinors in ref. VanHemelryck:2022ynr . It is given by
$P=\mathrm{e}^{\mathcal{K}/2}\mathcal{W}\,,$ (43)
where the functions $\mathcal{K}$ and $\mathcal{W}$ are determined by the fluxes and geometry:
$\mathcal{W}=-\frac{1}{4}\int\mathrm{e}^{-\phi}\left(\star\Phi\wedge F_{3}-F_{7}\right)+\frac{1}{8}\int\mathrm{e}^{-2\phi}\left(\Phi\wedge\mathrm{d}\Phi\right)\,,\quad\mathcal{K}=-4\log\left[\int\mathrm{e}^{-2\phi}\text{vol}_{7}\right]\,.$ (44)
The supersymmetry conditions (15)-(19) are equivalent to the condition that $\partial_{i}P=0$ for all eight scalar fields. Again, it is more straightforward to calculate the mass matrix by trading $\phi$ for $\delta$. The masses are usually expressed in AdS units, therefore it is useful to observe that $V=\Lambda=-2L_{\mathrm{AdS}}^{-2}$ and hence
$m^{2}L_{\mathrm{AdS}}^{2}=\mathrm{Eig}(M)\,,\qquad M^{ij}=\left(-\frac{2}{V}\right)\sqrt{g}^{ik}(\partial_{kl}V)\sqrt{g}^{lj}\,.$ (45)
The conformal dimensions of the field theory operators dual to the scalars are given by the simple relation
$\Delta=1+\sqrt{1+m^{2}L_{\mathrm{AdS}}^{2}}\,,\qquad m^{2}L_{\mathrm{AdS}}^{2}=\Delta(\Delta-2)\,.$ (46)
For the vacua on $\tilde{\mathbb{T}}^{7}/\hat{\Gamma}$, one gets the following masses and conformal dimensions:
$m^{2}L_{\mathrm{AdS}}^{2}$ =$=\left\{3,8,8,8,8,\left(6\frac{\omega_{a}}{\omega_{\Sigma}}\right)^{2}-1,\left(6\frac{\omega_{b}}{\omega_{\Sigma}}\right)^{2}-1,\left(6\frac{\omega_{c}}{\omega_{\Sigma}}\right)^{2}-1\right\}$ (47)
$\Delta$ =$=\left\{3,4,4,4,4,1+6\frac{|\omega_{a}|}{|\omega_{\Sigma}|},1+6\frac{|\omega_{b}|}{|\omega_{\Sigma}|},1+6\frac{|\omega_{c}|}{|\omega_{\Sigma}|}\right\}\,.$ (48)
So the conformal dimensions are all integer when $|\omega_{a,b,c}/\omega_{\Sigma}|$ are elements of $\mathbb{Z}/6$. This happens for the most symmetric case, $\omega_{a}=\omega_{b}=\omega_{c}$, but also for $-\omega_{a}=\omega_{b}=\omega_{c}$, for example. Note that for these cases, all masses are non-negative as well, and hence all geometric moduli are stabilised. The tadpole constraint only limits the size of $\omega_{\Sigma}=\omega_{a}+\omega_{b}+\omega_{c}$, still allowing for various spectra.
For the models on $\tilde{\mathbb{T}}^{7}/\check{\Gamma}$, the masses and corresponding conformal dimensions are:
$m^{2}L_{\mathrm{AdS}}^{2}=\{48,48,48,8,8,8,8,0\}\,,\qquad\Delta=\{8,8,8,4,4,4,4,2\}\,.$ (49)
Hence one of the scalars is massless but does not seem to correspond to a flat direction of the potential as the scalars are fixed, similar to refs. Becker:2024ayh ; Becker:2024ijy ; Rajaguru:2024emw . Notice that this spectrum is contained within the spectra of the non-supersymmetric scale-separated vacua of ref. Arboleya:2024vnp . That is quite surprising, as the authors there were looking at similar vacua with one O5-plane but with an SO(3)-invariant Ansatz that would require all structure constants $\omega_{a,b,c}$ to be the same (similar for all $f_{3,i}$’s). However, taking that choice here, but also artificially changing the fluxes $F_{246}\to-F_{246}$, $F_{235}=-F_{235}$, would amount to breaking supersymmetry, but corresponds to the solution found in ref. Arboleya:2024vnp . Applying these changes does result in the same expression for the scalar potential after carefully choosing $F_{136}=-F_{235}=F_{145}=-2F_{246}$ and $-\omega_{a}=\omega_{b}=\omega_{c}$ and hence can be understood as a “skew-whiffing” procedure Duff:1986hr ; Duff:1984sv , although breaking with the SO(3)-invariant Ansatz.
5 Summary
In this paper, supersymmetric AdS3 vacua of type IIB string theory with the parametric scale separation feature are constructed. These arise as compactifications on seven-dimensional solvmanifolds with co-closed $G_{2}$-structure, which can be seen as twisted seven-tori. The orbifolds and orientifold are carefully chosen such that all one- and two-forms are projected out and that, in one case, the geometry admits four sets of O5-planes and in another case only one. This means that some components of the $F_{3}$ Bianchi identity are not restricted and lead to unbounded fluxes. In principle, this could also be achieved by having seven sets of O5-planes as in ref. Emelin:2021gzx (so orbifolds without shifts), but it requires including D5-branes to cancel the O5-charge in some directions. It also leads to orbifold singularities for which it is unknown how to resolve them, even for the flat seven-torus. Additionally, such orbifolds also generate an O9-plane image and therefore the model resides in type I string theory instead (after adding 32 D9-branes). Working with the different orbifolds of this paper does not require the inclusion of D5-branes and keeps the setup in type IIB. Therefore, the solutions presented here are, as far as the author knows, the first supersymmetric, and parametrically scale-separated solutions of type IIB string theory at large volumes of the cycles and weak coupling, in the limit where the orientifold planes are smeared. Furthermore, the solution with only one set of O5-planes can be understood as a supersymmetric cousin of one of the vacua of ref. Arboleya:2024vnp through the skew-whiffing procedure. Moreover, the mass spectra of all the supersymmetric solutions, at tree level, lead to operators in the putative dual field theory with integer conformal dimensions, very similar to the investigations of refs. Conlon:2020wmc ; Apers:2022vfp ; Ning:2022zqx for 4d scale-separated vacua and contrary to the 3d scale-separated type IIA vacua, as explored in ref. Apers:2022zjx . The 3d solutions presented here share even more similarities with their 4d type IIA counterparts, especially the ones in massless type IIA: they rely on intersecting O-planes to stabilise the moduli and the internal manifold has non-trivial intrinsic torsion. The fact that the O-planes seem to intersect might be a worry as such setups are not well understood, also not at the world-sheet level. However, ref. Junghans:2023yue showed that for the 4d DGKT vacua DeWolfe:2005uu ; Camara:2005dc on the specific $\mathbb{T}^{6}/\mathbb{Z}_{3}^{2}$ orbifold, the four sets of O6-planes are not intersecting after blowing up the orbifold singularities. Although it is not entirely clear how the orbifold singularities of the twisted tori should be resolved, one could expect that a blow-up procedure, similarly for flat $\mathbb{T}^{7}$’s, might do so. The blow-up moduli could then be stabilised by balancing local RR three-form flux against the bulk seven-form flux, analogous to the blow-up procedure in DGKT DeWolfe:2005uu . A second reason for worry might be that the O-planes are smeared over the directions transverse to them, not being treated as the local sources that they are. Although computing the local O-plane backreaction for the 3d vacua is a computationally contrived problem here, previous successful attempts for other AdS vacua seem to indicate that there are no fundamental objections to it in principle. Indeed, in massive type IIA, refs. Baines:2020dmu ; Junghans:2020acz ; Marchesano:2020qvg ; Emelin:2022cac ; Andriot:2023fss ; Emelin:2024vug showed that on various backgrounds without intrinsic torsion, the O-plane backreaction is small and can be computed perturbatively, at first order. Additionally, refs. Cribiori:2021djm ; VanHemelryck:2024bas argued the same for backgrounds of massless type IIA with intrinsic torsion, where calculating the backreaction was essential to lift the solutions to M-theory. Finally, although the solutions in this paper rely on a specific class of seven-dimensional solvmanifolds, it would be interesting to explore other co-closed $G_{2}$-structures that yield similar scale-separated solutions in type IIB string theory. Indeed, the ideas presented here suggest that the search for new scale-separated vacua is not over yet.
Acknowledgements
I want to thank Adolfo Guarino, Matteo Morittu, Thomas Van Riet, Max Hübner, Magdalena Larfors and especially Álvaro Arboleya and George Tringas for interesting discussions. I am financially supported by the Olle Engkvists Stifelese.